OpenStudy (anonymous):

Which of these is a literal equation? ten minus six equals four three squared plus four equals ten plus three x squared minus seven equals y plus four three time eight minus ten equals three time negative two

OpenStudy (nurali):

they are all literal since you wrote them down x squared minus seven equals y plus four is quadratic since it has a squared term three time eight minus ten equals three time negative two can be either 3 *(8-10) = 3 * -2 or (3*8)-10 = 3*2 one is true one is not.

OpenStudy (texaschic101):

I am guessing the third one. A literal equation is an equation that contains two or more variables. In the third answer choice, you have variable x and variable y.